OpenStudy (anonymous):

Use a table to find two consecutive integers between which the solution lies. 4.5=1.4-0.4y

OpenStudy (anonymous):

1.4-0.4y=4.5 1.4-4.5=0.4y -3.1/0.4=0.4y/0.4 -7.75=y Check my math